C=200+0.75Yd G=200 I=100 T=100
Derive savings function
Calculate equilibrium level.
Compute multipliers of investment and taxes.
Suppose investment rises by 100,find new equilibrium.
Suppose government reduces expenditure by 100,how much would national income rise?
If desired level of income is 2400, what level of government spending would be set?
1
Expert's answer
2020-09-09T11:02:00-0400

Y=C+I+G

Yd=Y-T

Y=200+0.75(Y-100)+100+200

Y-0.75Y=425

0.25Y=425

Y=1700

C=200+0.75(1700-100)=1400

Y=C+S S=Y-C=1700-1400=300

"mi=\frac{1}{1-MPC}=\frac{1}{1-0.75}=4"

"mt=\frac{-MPC}{1-MPC}=\frac{-0.75}{1-0.75}=-3"

I=200

Y=200+0.75(Y-100)+200+200

Y-0.75Y=525

0.25Y=525

Y=2100


G=100

Y=200+0.75(Y-100)+100+100

Y-0.75Y=325

0.25Y=325

Y=1300


Y=2400


Y=200+0.75(Y-100)+100+G

2400=200+0.75(2400-100)+100+G

2400=200+1725+100+G

G=375